I need examples of when the sum of a positive number and several negative numbers will be positive.

Found 2 solutions by rfer, RAY100:
Answer by rfer(16322)

As long as the positive number is larger than the sum of the negative numbers, the answer is positive.
Example:
+6-2-3=+1

Answer by RAY100(1637)

Let P = positive number,,,and, N= sum of negative numbers
.
Given,,P+N>0
.
P>-N, that is,,Positive Number greater than sum of negative numbers
.
Examples, P=5,,N=4=1+1+1+1, or 5-1-1-1-1>0
.
or,,,,,P=5, N=3=1+1+1,,,,or 5-1-1-1>0
.
or,,,,,,P=5, N=2=1+1,,,,,or 5-1-1 >0